help me solve: The sum of 2 numbers is 40 and their difference is 18
Answer by nerdybill(5411)

The sum of 2 numbers is 40 and their difference is 18
.
Let x = one of two numbers
and y = the second number
.
x + y = 40
x - y = 18
.
Using the "elimination method" add the two equations together:
x + y = 40
x - y = 18
-----------
2x = 58
x = 29
.
Substitute the above into:
x + y = 40
29 + y = 40
y = 40 - 29
y = 11
.
The two numbers are 29 and 11